Tracy can wallpaper 2 rooms in a new house in 10 hours. Together with her trainee they can wallpaper the 2 rooms in 7 hours. How long would it take the trainee working by herself to do the job?
Answer by richwmiller(17219)

In one hour Tracy can do 1/10 of the job
x/10+x/y=1
x=7
y=23 1/3 hours for the trainee
check
x/10+x/70/3=1
x/10+3x/70=1
7x/70+3x/70=1
10x=70
x=7
ok